OpenStudy (anonymous):

What is the inverse of the function f(x) = x +3? a. h(x) = x + 3 b. h(x) = x – 3 c. h(x) = x – 3 d, h(x) = x + 3

OpenStudy (anonymous):

@madhu.mukherjee.946

OpenStudy (anonymous):

wait i wrote it wrong @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

okay

Directrix (directrix):

These two options appear to be identical: b. h(x) = x – 3 c. h(x) = x – 3 a and d also appear the same.

Directrix (directrix):

@madhu.mukherjee.946 Please show the work on how you got that answer. Thanks.

OpenStudy (madhu.mukherjee.946):

@Directrix okay

OpenStudy (anonymous):

a. h(x) = 1/3x + 3 b. h(x) = 1/3x – 3 c. h(x) = x – 3 d. h(x) = x + 3